# Resource Public Key Infrastructure BGP Border Gateway Protocol

### v1

@@ IV-C MFT and CRL
Every CA in RPKI requires a signed manifest and a signed CRL to ensure integrity and track revocation of repository content. This sets RPKI apart from other technologies like HTTPS, which only includes CRLs and therefore has smaller per-CA overhead (1 object instead of 2), raising the question why RPKI increases complexity by using two such objects. RFC6486 states that manifests are modeled after CRLs, sharing many characteristics, like their fundamental structure and their use of serial numbers to assure a consistent sequence between new issuance of CRL / MFT respectively. Further, both serve similar roles in the architecture: Ensuring validity of objects from the perspective of the CA. Their combined role is acknowledged in the RFCs, stating that manifests and CRLs should always be re-issued together. This advice is generally followed by operators, and we find all production CAs re-issue their CRL if they re-issue the manifest, and vice versa. Any change in the repository, like a new or deleted object or a refresh of validity will thus always lead to re-issuance of manifest and CRL. Despite their similar function and even though they always must be changed together, manifest and CRLs are separate, individually signed objects. Keeping the objects separate creates overhead for issuance, downloading, cache size, and validation effort.
@@ IV-D RRDP File Structure
RPKI uses RRDP as its primary data exchange protocol, which utilizes XML-encoding for all files. The RFC does not give reasoning why XML was chosen, however, using XML likely follows the design philosophy of using existing and accepted technologies in RPKI (RFC6480).
@@ V-A Bandwidth
Currently, 91 repositories and 4211 RPs are deployed in production RPKI systems, measured on April 17 2025. Each of these repository servers must offer sufficient bandwidth to accommodate requests by all RPs, which regularly poll all servers. The exact required bandwidth is a function of the number of RPs that download the data, the size of the files within the repository, and whether RPs download deltas or snapshots. In normal operation, most RPs download deltas, which are generally smaller than snapshots (RFC8182). But cold starts, session refreshes, or faulty deltas lead to downloading complete snapshots, and the repository must ensure bandwidth is sufficient for snapshot downloads by all RPs [^4] Link IETF.
Mapping the current global RPKI state, the average RRDP snapshot size is around 15 MB. Combining this size with data on RP distribution [^5] https://rov-measurements.nlnetlabs.net/stats/ (Accessed April 17 2025), we can approximate the required minimal bandwidth of repositories to around 670 Mbit/s. An interested reader is referred to Appendix -A for detailed computations. This bandwidth estimate provides a lower bound, not accounting for HTTPS overhead, retransmissions, and RP fallbacks. In practice, repositories need to provide more than 670 Mbit/s bandwidth to ensure sufficient resilience against network errors and attacks, requiring repositories to host their servers on high-availability, high bandwidth infrastructure.
Since RPKI is continuously developing and new use-cases are still being standardized, the amount of RPs and the amount of data in repositories is growing [^6] Link IETF. Currently, ca. 55% of globally announced address space is covered by RPKI ROAs and around 20% of users are protected by ROV enforcement using an RP [9, 7]. Extrapolating RPKI deployment to 100% ROA and 100% ROV coverage yields a doubling of average snapshot size to 30MB, and a 5x increase in the amount of RPs to 21,055. Thus, in a conservative approximation on full RPKI deployment average repositories will require minimal bandwidth of 6.7 GBit/s. For larger repositories, the required data-rate is even larger. Applying the same approximation to the RIPE NCC snapshot, which is currently 230MB, RIPE will have to provide at least 102 GBit/s bandwidth. The bandwidth requirements of current RPKI thus do not scale well for growing adoption of RPKI, even under best-case assumptions on use-cases, congestion, and other networking errors. More objects and snapshots additionally increase the fetch durations of RPs, exasperating issues with low flexibility and resilience.
@@ V-B High Latency Fetch Intervals
Setting up current RPs and running a full fetch in a default configuration takes 7min on Routinator, 7.1min on rpki-client, and 10min in Fort, with between 71% (Routinator) and 58% (Fort) of time spent on downloading RPKI data, and 29% (42%) of time spent on processing / validation of data [^7] We run all RPs sequentially on a machine with 100 Mbit/s available bandwidth and 8-core i7 processor.. Investigating the RPs, we find that if fetch times increase beyond the thresholds of RPs, it will lead to data loss; we provide more detailed considerations on threshold limits and object expiration in Appendix -B. Full RPKI adoption will increase RP fetch times, exceeding the thresholds.
Furthermore, fetches include downloading of around 1.2 GB of data, and requires validation of around 764000 RSA signatures, also contributing to high latency fetch intervals if resources are insufficient [^8] Link IETF. To illustrate how constrained environments can impact RPs, we repeat the RP measurement with bandwidth constrained to 10 Mbit/s. This increases total validation time 3x, with Routinator and rpki-client taking around 21min to finish and Fort taking 24min. We provide more detailed considerations on the implications of long fetches in Appendix -C.
@@ V-C Insecurity Through Complexity
Looking at all CVEs for all actively maintained RPs and all RPKI papers published in the last five years, we find a total of 47 vulnerabilities. The majority (29/47) of these vulnerabilities stem from errors in the decoding and processing of RPKI objects. This is not surprising, as the decoding and processing of RPKI objects incurs substantial complexity, including DER decoding of at least six different currently standardized object types, and each RPKI object including interconnected validation steps. 7 vulnerabilities in RPKI were discovered in the RPKI-specific data exchange protocol RRDP, including issues with compression, XML-specific attacks, and content decoding, and path-traversal [21, 15, 13].
@@ V-D Algorithm Agility in RPKI
Currently, RPKI supports solely RSA-2048 for its public key signatures (RFC6485), but the need to transition to new algorithms is acknowledged in RFC6916. This transition will be required for supporting quantum-secure algorithms, a shift the security community aims to complete roughly until 2035  [^9] Document EU, Document UK, Document USA. Depending on which algorithm will be implemented into RPKI, the overall size of the downloaded RPKI data will increase from today 1.2 GB to between 2.9 GB and 39.1 GB. For more details on size computations, see Appendix -D. As all proposed post-quantum algorithms increase overall object size, transitioning to a post-quantum secure algorithm will thus put additional load on the RPKI architecture.
@@ VI-A Design Goals
Our design goals ensure that the improvements to RPKI are secure, deployable, and maintainable.
@@
#3 Re-use existing solutions. RPKI heavily utilizes existing concepts and solutions, like ASN.1 for encoding and X.509 for certificates, a design goal specifically mentioned in RFC6480. This is sensible to ensure minimal development overhead for implementations and improve acceptance by the community. Any improvements to the design should follow the same principle. Where possible, the improved design re-uses the current RPKI structure, or other established technologies, and keeps changes as minimal as possible.
#4 Major Improvements. Since RPKI is already deployed, there is a natural aversion against changes to the protocol that may cause disruptions or other issues in production systems. Thus, all proposed adaptions must provide major and tangible improvements to RPKI operators that warrant the inherent risk of changes to production systems. Specifically, improvements should significantly lower processing times, download times, or reduce memory footprint. To quantify what constitutes a major change, we run our own RPKI RP software and observe natural fluctuations of the above described metrics. We find natural fluctuations of up to 5% between consecutive runs of the RPs [^10] For download times, we see more significant fluctuations of up to 50% due to changes in repository reachability. On consecutive runs without changes in reachability of remote repositories, the 5% observation holds.. We define an improvement as significant if it improves one of the above metric by more than double its natural fluctuations, i.e., improves by at least 10%. All changes we make in iRPKI fulfill this requirement.
@@ VI-B Removing the EE-Certificate
Figure 2: ROA structure.
Do EE-certificates improve security? Using EE-certificates in RPKI does not provide stronger security than signing the object with the CA key, as both use the same signing procedure and cryptographic cipher. The support for built-in EE-certificates is motivated by (validation-) convenience and object self-sufficiency, not added security (RFC5652). Adapting the design to change or remove the EE-certificate is, however, not trivial. The current RPKI design relies on EE-certificates for validation, for example to contain the signing key and validity period of an object. Further, EE-certificates also contain multiple fields essential for RP processing, like the parent key fingerprint or owner name. Making changes to EE-certificates thus requires careful considerations on how each EE-certificate field is used in object validation.
@@
Delegations. EE-certificates contain the IpAddrDelegation and ASNumberDelegation extensions to attest resources allocated to this certificate. The validation of the IP and ASN resources is conducted through the CA certificate. The RFC does not make clear why these delegations are included in EE-certificates, and the manifest RFC6486 even mandates that both extensions are set to null (inherit). ROAs need to have explicit values in the extension according to the RFC and RPs must validate them. However, ROA validation does not make use of the values in the ASN delegation, and while it validates that the ROA resources are contained in the IP resources of the EE-certificate, these resources must be a subset of the CA resources anyway, so there is no clear benefit of additionally including IP address delegation validation in the EE-certificate, after it was already validated through the CA certificate. Our tests show that both fields could be omitted or set to inherit in all signed objects without reducing security / authenticity, as validation of all RPs will fail if these resources conflict with the resources stated in the CA certificate.
@@
Parsing signed objects without a certificate is already supported by existing software, like OpenSSL, as included certificates are optional as per CMS RFC5652. Investigating the implementations, we find that Routinator has a strict custom implementation of signed object parsing, and requires minor code adaptions ($\approx$ 20 Lines of Code) to remove the need for the existence of an EE-certificate in parsing.
@@
Signatures on RPKI objects ensure authenticity, as RPs can validate the signature of the CA to confirm that the CA has authorized the signed object. Signatures do not ensure integrity: A malicious repository may still delete some objects issued by the CA, compromising integrity and availability. To detect such attacks, RPKI uses the manifest file. RPs compare downloaded repository content against the manifest and fail validation of the entire CA if manifest validation fails, as the failed validation indicates an attack on repository integrity, like deletion of an object. The manifest, however, not only detects integrity compromise but additionally offers a signed protection of authenticity: If an attacker changes any object covered by the manifest, like a ROA, the hash of the tampered file will differ from the hash in the manifest, leading manifest validation to fail. Since manifest validation failure invalidates the entire CA, this prevent acceptance of tampered objects. Since the manifest is signed by the CA, the attacker can not manipulate the manifest hash to the new malicious object content to circumvent validation failure. While the initial design goal of manifests was integrity, they thus offer a redundant mechanism to verify object authenticity. In practice, if an attacker manipulates a ROA, breaking authenticity, the validation will not fail when the RP validates the ROA signature, but when the RP validates the hash of this ROA in the manifest. ROA signature validation and manifest hash list validation both detect compromised object authenticity.
@@
This authenticity protection of manifests is possible in a threat model of an external attacker, i.e., an attacker without ownership of the private key (RFC6486). If the attacker has ownership of the CA private key, they can adapt the ROA hash on the manifest content. However, in this case, the ROA signature would also not protect the CA, as the attacker can create valid ROA signatures using the CA key.
@@
In the current design of RPKI, signatures on ROAs merely add an additional means to detect tampering of the object, but do not improve overall security. ROA authenticity is validated through the manifest, and signature validation of ROAs will only fail for mistakes by the CA, not in malicious attacks. At the same time, since ROA are the most abundant objects in the RPKI (76% of all RPKI objects are ROAs), validating their signatures takes up a major part of total RP processing time. Removing signature from ROAs does not reduce security while providing substantial benefits in lower validation load on RPs, and smaller size of ROA files.
@@
ROA without signature. Removing ROA signatures does not require adaptions of manifest content, but modifies its stated role in RFC9286 from integrity protection to additional authenticity protection. This has implications on the cryptographic and availability requirements for the manifest. Currently, manifests use sha256 for hash validation, which is considered secure for cryptographic applications [3]. Removing ROA signatures requires strict availability of manifests, as ROAs can not be validated without it. This requirement is already part of current RPKI, and all RPs will fail validation of any repository if the manifest is unavailable. RPs will not accept any ROAs that are not listed on the manifest or have invalid manifest hashes, even if they are validly signed. The new role of the manifest thus does not lead to any changes in threat model or availability requirements, and RPs do not need to implement any new checks. We discuss the implications of removing ROA signatures for potential future use-cases of RPKI in Section VIII.
@@ VI-F Improved RRDP File Structure and Encoding
Protobuf over XML. Even with the improvements, XML elements still result in overhead. Human readability of XML structure elements, like <<<snapshot>>>, take up significant space, and the requirements for XML parsing and validation create substantial processing effort in the RPs (11.5% of total processing time in Routinator). Using a different encoding format than XML improves size and processing speed of RRDP files. Looking at previous research [4, 19], we find protobuf as the most promising solution to improve RRDP files, as it enables smaller file sizes and more efficient parsing than XML. Using our above derived optimizations, we develop protobuf formats for RRDP notification, Snapshot, and Delta respectively. All designs are derived from the existing improved XML structure, translated into protobuf templates. We provide protobuf definitions for all RRDP files in Appendix B-A.
@@
The downside of protobuf compared to ASN.1 is reduced flexibility. While ASN.1 generally allows for extensions, a wide range of types and flexible encoding, protobuf requires strict object definitions, has limited types and only supports one specific encoding. However, this strictness of protobuf is not a limitation in the context of RPKI. The RPKI RFCs specifically express that RPKI is designed to be strict and tailored to its very specific use-case (RFC6487). The RFCs only allow specific types, one encoding, and disallow any extensions or optional fields. This ensures deterministic structure of RPKI objects, preventing real-world object structure to diverge overtime. The strictness also omits the benefits of using a technology like ASN.1, that supports many other use-cases and provides templates for other object types, which are neither necessary nor desired by the RFCs in the context of RPKI. We provide protobuf definitions for improved ROAs and improved manifests in Appendix B-B and B-C.

### v2

@@ IV-C MFT and CRL
Every CA in RPKI requires a signed manifest and a signed CRL to ensure integrity and track revocation of repository content. This sets RPKI apart from other technologies like HTTPS, which only includes CRLs and therefore has smaller per-CA overhead (1 object instead of 2), raising the question why RPKI increases complexity by using two such objects. [1] states that manifests are modeled after CRLs, sharing many characteristics, like their fundamental structure and their use of serial numbers to assure a consistent sequence between new issuance of CRL / MFT respectively. Further, both serve similar roles in the architecture: Ensuring validity of objects from the perspective of the CA. Their combined role is acknowledged in the RFCs, stating that manifests and CRLs should always be re-issued together. This advice is generally followed by operators, and we find all production CAs re-issue their CRL if they re-issue the manifest, and vice versa. Any change in the repository, like a new or deleted object or a refresh of validity will thus always lead to re-issuance of manifest and CRL. Despite their similar function and even though they always must be changed together, manifest and CRLs are separate, individually signed objects. Keeping the objects separate creates overhead for issuance, downloading, cache size, and validation effort.
@@ IV-D RRDP File Structure
RPKI uses RRDP as its primary data exchange protocol, which utilizes XML-encoding for all files. The RFC does not give reasoning why XML was chosen, however, using XML likely follows the design philosophy of using existing and accepted technologies in RPKI [20].
@@ V-A Bandwidth
Currently, 91 repositories and 4211 RPs are deployed in production RPKI systems, measured on April 17 2025. Each of these repository servers must offer sufficient bandwidth to accommodate requests by all RPs, which regularly poll all servers. The exact required bandwidth is a function of the number of RPs that download the data, the size of the files within the repository, and whether RPs download deltas or snapshots. In normal operation, most RPs download deltas, which are generally smaller than snapshots [4]. But cold starts, session refreshes, or faulty deltas lead to downloading complete snapshots, and the repository must ensure bandwidth is sufficient for snapshot downloads by all RPs [3].
Mapping the current global RPKI state, the average RRDP snapshot size is around 15 MB. Combining this size with data on RP distribution [31], we can approximate the required minimal bandwidth of repositories to around 670 Mbit/s. An interested reader is referred to Appendix -A for detailed computations. This bandwidth estimate provides a lower bound, not accounting for HTTPS overhead, retransmissions, and RP fallbacks. In practice, repositories need to provide more than 670 Mbit/s bandwidth to ensure sufficient resilience against network errors and attacks, requiring repositories to host their servers on high-availability, high bandwidth infrastructure.
Since RPKI is continuously developing and new use-cases are still being standardized, the amount of RPs and the amount of data in repositories is growing [34]. Currently, ca. 55% of globally announced address space is covered by RPKI ROAs and measurements using representative samples show around 28% of ASes and 20% of users are protected by ROV enforcement using an RP [21, 12, 17]. Extrapolating RPKI deployment to 100% ROA and 100% ROV coverage - assuming linear ROA growth and a stable ratio of RPs per enforcing AS - yields a doubling of average snapshot size to 30MB, and a 5x increase in the amount of RPs to 21,055. Thus, in a conservative approximation on full RPKI deployment average repositories will require minimal bandwidth of 6.7 GBit/s. For larger repositories, the required data-rate is even worse. Applying the same approximation to the RIPE NCC snapshot, which is currently 230MB, RIPE will have to provide at least 102 GBit/s bandwidth. The bandwidth requirements of current RPKI thus do not scale well for growing adoption of RPKI, even under best-case assumptions on use-cases, congestion, and other networking errors. More objects and snapshots additionally increase the fetch durations of RPs, exasperating issues with low flexibility and resilience. While longer default RP fetch intervals could lighten load on repositories, long intervals are disadvantages since they extend propagation delay between the upload of new objects and impact in routing. For example, if a ROA is uploaded but fetch intervals are long, new BGP announcements can conflict the old ROAs in all RPs that have not yet updated their cache, leading to invalidation of BGP announcements and thereby loss of traffic. The longest current default interval of production RPs is 1h.
@@ V-B High Latency Fetch Intervals
Setting up current RPs and running a full fetch in a default configuration takes 7min on Routinator, 7.1min on rpki-client, and 10min in Fort, with between 71% (Routinator) and 58% (Fort) of time spent on downloading RPKI data, and 29% (42%) of time spent on processing / validation of data [^2] We run all RPs sequentially on a machine with 100 Mbit/s available bandwidth and 8-core i7 processor.. Investigating the RPs, we find that if fetch times increase beyond download thresholds, it will lead to data loss; we provide more detailed considerations on threshold limits and object expiration in Appendix -B. Full RPKI adoption will increase RP fetch times, increasing risks of exceeding the thresholds. Contributing to this issue, the number of repositories is continuously growing [11], which is extending RP fetch times further. Multiple previous research works raised issues with scalability of the architecture with growing numbers of repositories [37, 11], as RPs have to contact each repository individually, and configuration errors in the growing number of repositories further extend fetches.
Furthermore, fetches include downloading of around 1.2 GB of data, and requires validation of around 764000 RSA signatures, also contributing to high latency fetch intervals if resources are insufficient [36]. To illustrate how constrained environments can impact RPs, we repeat the RP measurement with bandwidth constrained to 10 Mbit/s. This increases total validation time 3x, with Routinator and rpki-client taking around 21min to finish and Fort taking 24min. We provide more detailed considerations on the implications of long fetches in Appendix -C.
@@ V-C Insecurity Through Complexity
Looking at all CVEs for the actively maintained RPs and all RPKI papers published in the last five years, we find a total of 47 vulnerabilities. The majority (29/47) of these vulnerabilities stem from errors in the decoding and processing of RPKI objects. This is not surprising, as the decoding and processing of RPKI objects incurs substantial complexity, including DER decoding of at least six different currently standardized object types, and each RPKI object including interconnected validation steps. 7 vulnerabilities in RPKI were discovered in the RPKI-specific data exchange protocol RRDP, including issues with compression, XML-specific attacks, content decoding, and path-traversal [40, 27, 25].
@@ V-D Algorithm Agility in RPKI
Currently, RPKI solely supports RSA-2048 for its public key signatures [16], but the need to transition to new algorithms is acknowledged in RFC6916. This transition will be required for supporting quantum-secure algorithms, a shift the security community aims to complete roughly until 2035 333www.europol.europa.eu/, www.ncsc.gov.uk/, nvlpubs.nist.gov/. Depending on which algorithm will be implemented into RPKI, the overall size of the downloaded RPKI data will increase from today 1.2 GB to between 2.9 GB and 39.1 GB. For more details on size computations, see Appendix -D. As all proposed post-quantum algorithms increase overall object size, transitioning to a post-quantum secure algorithm will thus put additional load on the RPKI architecture.
@@ VI-A Design Goals
Our conceptual design goals ensure that the improvements to RPKI are secure, deployable, and maintainable.
@@
#3 Re-use existing solutions. RPKI heavily utilizes existing concepts and solutions, like ASN.1 for encoding and X.509 for certificates, a design goal specifically mentioned in [20]. This is sensible to ensure minimal development overhead for implementations and improve acceptance by the community. However, re-using existing solutions incurs a trade-of between resulting overhead and gained benefits. Any changes to the design should thus weigh benefits against downsides. For this, we evaluate each proposed change for its complexity and performance impact. Where changes do not incur substantial benefit, the improved design re-uses the current RPKI structure, and generally keeps changes as minimal as possible.
#4 Major Improvements. Since RPKI is already deployed, there is a natural aversion against changes to the protocol that may cause disruptions or other issues in production systems. Thus, all proposed adaptions must provide major and tangible improvements to RPKI operators to warrant the inherent risk of changes to production systems. Specifically, improvements should significantly lower processing times, download times, or reduce memory footprint. To quantify what constitutes a major change, we run our own RPKI RP software and observe natural fluctuations of the above described metrics. We find natural fluctuations of up to 5% between consecutive runs of the RPs [^4] Download times fluctuate up to 50% due to changes in repository reachability. On consecutive runs without changes in reachability of remote repositories, the 5% observation holds.. We define an improvement as significant if it improves one of the above metric by more than double its natural fluctuations, i.e., improves by at least 10%. All changes we make in iRPKI fulfill this requirement.
@@ VI-B Removing the EE-Certificate
Figure 2: ROA object structure in current RPKI.
Do EE-certificates improve security? Using EE-certificates in RPKI does not provide stronger security than signing the object with the CA key, as both use the same signing procedure and cryptographic cipher. The support for built-in EE-certificates is motivated by (validation-) convenience and object self-sufficiency, not added security [15]. Adapting the design to change or remove the EE-certificate is, however, not trivial. The current RPKI design relies on EE-certificates for validation, for example to contain the signing key and validity period of an object. Further, EE-certificates also contain multiple fields essential for RP processing, like the parent key fingerprint or owner name. Making changes to EE-certificates requires careful considerations on how each EE-certificate field is used in object validation.
@@
Delegations. EE-certificates contain the IpAddrDelegation and ASNumberDelegation extensions to attest resources allocated to this certificate. The validation of the IP and ASN resources is conducted through the CA certificate. The RFC does not make clear why these delegations are included in EE-certificates and in manifests, both extensions must be set to null (inherit) [1]. ROAs need to have explicit values in the extension according to the RFC and RPs must validate them. However, ROA validation does not make use of the values in the ASN delegation, and while it validates that the ROA resources are contained in the IP resources of the EE-certificate, these resources must be a subset of the CA resources anyway, so there is no clear benefit of additionally including IP address delegation validation in the EE-certificate, after it was already validated through the CA certificate. Our tests show that both fields could be omitted or set to inherit in all signed objects without reducing security / authenticity, as validation of all RPs will fail if these resources conflict with the resources stated in the CA certificate.
@@
Parsing signed objects without a certificate is already supported by existing software, like OpenSSL, as included certificates are optional as per CMS [15]. Investigating the implementations, we find that Routinator has a strict custom implementation of signed object parsing, and requires minor code adaptions ($\approx$ 20 Lines of Code) to remove the need for the existence of an EE-certificate in parsing.
@@
Signatures on RPKI objects ensure authenticity, as RPs can validate the signature of the CA to confirm that the signed object was validly issued. Signatures do not protect integrity: A malicious repository may still delete some objects issued by the CA, compromising integrity and availability. To detect such attacks, RPKI uses the manifest file. RPs compare downloaded repository content against the manifest and fail validation of the entire CA if manifest validation fails, as the failed validation indicates an attack on repository integrity, like deletion of an object. The manifest, however, not only detects integrity compromise but additionally offers a signed protection of authenticity: If an attacker changes any object covered by the manifest, like a ROA, the hash of the tampered file will differ from the hash in the manifest. This leads manifest validation to fail, invalidating the entire CA and preventing acceptance of tampered objects. Since the manifest is signed by the CA, the attacker can not manipulate the manifest hash to the new malicious object content to circumvent validation failure. While the initial design goal of manifests was integrity, they thus offer a redundant mechanism to verify object authenticity. In practice, if an attacker manipulates a ROA, breaking authenticity, the validation will not fail when the RP validates the ROA signature, but when the RP validates its manifest hash. ROA signature validation and manifest hashes both detect compromised object authenticity.
@@
This authenticity protection of manifests is possible in a threat model of an external attacker, i.e., an attacker without ownership of the private key [1]. If the attacker has ownership of the CA private key, they can adapt the ROA hash on the manifest content. However, in this case, the ROA signature would also not protect the CA, as the attacker can forge valid ROA signatures using the CA key.
@@
In the current design of RPKI, signatures on ROAs merely add an additional means to detect tampering of the object, but do not improve overall security. ROA authenticity is validated through the manifest, and signature validation of ROAs will only fail for mistakes by the CA, not in malicious attacks. At the same time, since ROA are the most abundant objects in the RPKI (76% of all RPKI objects are ROAs), validating their signatures takes up a major part of total RP processing time. Removing signatures from ROAs does not reduce security while providing substantial benefits in lower validation load on RPs, and smaller size of ROA files.
@@
ROA without signature. Removing ROA signatures does not require adaptions of manifest content, but modifies its stated role in [1] from integrity protection to additional authenticity protection. This has implications on the cryptographic and availability requirements for the manifest. Currently, manifests use sha256 for hash validation, which is considered secure for cryptographic applications [8]. Removing ROA signatures requires strict availability of manifests, as ROAs can not be validated without it. This requirement is already part of current RPKI, and all RPs will fail validation of any repository if the manifest is unavailable. RPs will not accept any ROAs that are not listed on the manifest or have invalid manifest hashes, even if they are validly signed. The new role of the manifest thus does not lead to any changes in threat model or availability requirements, and RPs do not need to implement any new checks. We discuss the implications of removing ROA signatures for potential future use-cases of RPKI in Section VIII.
@@ VI-F Improved RRDP File Structure and Encoding
Protobuf over XML. Even with the improvements, XML elements still result in overhead. Human readability of XML structure elements, like <<<snapshot>>>, take up significant space, and the requirements for XML parsing and validation create substantial processing effort in the RPs (11.5% of total processing time in Routinator). Using a different encoding format than XML improves size and processing speed of RRDP files. Looking at previous research [9, 38], we find protobuf as the most promising solution to improve RRDP files, as it enables smaller file sizes and more efficient parsing than XML. Using our above derived optimizations, we develop protobuf formats for all RRDP files, derived from the existing improved XML structure. We provide protobuf definitions for all RRDP files in Appendix B-A. Using protobuf additionally removes susceptibility of RPs to XML-based attacks, like XML-bombs [35], but makes them potentially vulnerable to attacks on protobuf parsing. Keeping software up-to-date to prevent attacks on parsing thus remains important also in iRPKI.
@@
The downside of protobuf compared to ASN.1 is reduced flexibility. While ASN.1 generally allows for extensions, a wide range of types and flexible encoding, protobuf requires strict object definitions, has limited types and only supports one specific encoding. However, this strictness of protobuf is not a limitation in the context of RPKI. The RPKI RFCs specifically express that RPKI is designed to be strict and tailored to its very specific use-case [18]. The RFCs only allow specific types, one encoding, and disallow any extensions or optional fields. This ensures deterministic structure of RPKI objects, preventing real-world object structure to diverge overtime. The strictness also omits the benefits of using a technology like ASN.1, that supports many other use-cases and provides templates for other object types, which are neither necessary nor desired by the RFCs in the context of RPKI. We provide protobuf definitions for improved ROAs and improved manifests in Appendix B-B and B-C.